5 White Socks and 3 Black socks are in a drawer. I take out two socks. What is the probability the socks are the same colour?

Problem Tackles: Counting Probabilities, Non-replacement probability problems. Keeping track of events
Total number of socks: 5 + 3 = 8Scenarios: Could pick two white socks or two black socksP(Picking two white Socks) = 5/8 * 4/7 = 20/56 . (Number of socks after first pick is 7, number of white socks left is 4)P (Picking two Black socks) = 3/8 *2/7 = 6/56 (same idea as above)P(picking two socks of same colour) = P(Picking two white socks) + P(Picking two black socks) = 6/56 + 20/56 = 26/56

BS

Related Maths GCSE answers

All answers ▸

How do I solve simultaneous equations?


Starting with x^2+2x+1=0 use the method of factorising to solve for x.


Solve x^2 - 5x - 14 = 0


3x + 12 = 24, solve for x.